# Magnetic phase diagram of multiferroic and magnetocaloric $\mathbf{TmFeO_{3}}$

### v2

@@ III.2 Magnetic scattering dependence on magnetic field
Figure 3: Schematic of the magnetic phase diagram for $\rm{TmFeO_{3}}$ a) external magnetic field along the b-axis b) along c-axis. Squares, circles and triangles are the inflection points of the temperature dependence function of the intensity for different reflections. Roman numbers designate regions corresponding by different magnetic representation: I – $\Gamma 4$; II – $\Gamma 24$; III – $\Gamma 2$; IV – $\Gamma^{\rm{Fe}}_{4}+\Gamma^{\rm{Tm}}_{2}$; V – $\Gamma^{\rm{Fe}}_{4}-\Gamma^{\rm{Tm}}_{2}$.
@@
Regarding temperature region 2 – 30 K and field range 2 – 4 T, one can suppose, that in this case external magnetic field favors for the decoupling between $\rm{Tm^{3+}}$ and $\rm{Fe^{3+}}$ ions and Fe subsystem recover high-temperature configuration $\Gamma 4$, while Tm subsystem remains in $\Gamma 2$. At higher fields like 5.8 T Tm subsystem also goes to $\Gamma 4$ configuration but with ferromagnetic component directed opposite to $\rm{Fe^{3+}}$ one. Authors of [44] observed a magnetic-field-induced spin reorientation $\Gamma 2$rightarrow $\Gamma 4$ at 1.6 K in a b-cut $\rm{TmFeO_{3}}$ single crystal in the static magnetic field range 2.2 – 3.6 T applied along b-axis using terahertz spectroscopy. We have also observed change of intensity of peaks at magnetic field applied along b-axis in the close field range that confirms phase transition. Situation with reverse reorientation transition without external magnetic field was observed in $\rm{TbFeO_{3}}$, a unique compound in which two types of spin reorientation from $\Gamma 4$ to $\Gamma 2$ at 8.5 K and then $\Gamma 2$ to $\Gamma 4$ at 3 K were reported [45].
Figure 4: Hysteresis of G-type peak ($-3,1,0$) intensity in magnetic fields along the b-axis at 20 K. Error bars are of the size of the symbols.
In addition, hysteresis of some peaks for the field along the b-axis was observed in the experiment, which is absent for the peaks with the field along the c-axis, i.e. along the giant magnetocaloric axis. The biggest hysteresis was observed for G-type peak $(-3,1,0)$, see Fig. 4, and a smaller hysteresis was observed for G$(1,1,0)$, G$(-1,1,0)$, A$(-4,1,0)$, A$(-2,1,0)$ and F$(-2,2,0)$ peaks (not shown). Temperature magnetization hysteresis was described using model with a few domains, which explain jump/drop of magnetization in the spin reorientation temperatures range in $\rm{TmFeO_{3}}$ [31].
@@ IV Conclusions
X-ray studies demonstrate that the lattice parameters and atomic coordinates change almost linearly in $\rm{TmFeO_{3}}$ through the spin reorientation transition, keeping the orthorhombic symmetry group Pnma. The direction of the electric dipole moments does not change at any magnetic phase. This is additional evidence that electric system of $\rm{TmFeO_{3}}$ does not have influence on the spin-reorientation transition in this compound. The main reason for the spin reorientation transition connected with superexchange interactions 4f-3d, or with more complicate scheme of magnetic interactions [40]. Application of external magnetic field influences significantly on the spin-reorientation processes, moving the temperatures of the transitions up or low, depending on the direction of the external field. Thus, the application of the field along the c axis creates a situation favorable for the existence of the $\Gamma 2$ configuration, which leads to an increase in temperature of transition from $\Gamma 4$ to $\Gamma 2$ phase. The field along b-axis of $\rm{TmFeO_{3}}$ favors for $\Gamma 4$ configuration. This result in a decrease in the transition temperature from the $\Gamma 4$ phase to the $\Gamma 2$ phase, an increase in the temperature range in which the mixed phase is observed; and also allows for the formation of new phases. At low temperatures below 20 K and fields between 2 – 4 T a new phase was observed, which has configuration supposedly $\Gamma 4$ for $\rm{Fe^{3+}}$ ions and $\Gamma 2$ for $\rm{Tm^{3+}}$. At higher fields at these low temperatures another phase emerges with configuration $\Gamma 4$ for both $\rm{Fe^{3+}}$ subsystem and $\Gamma 4$ for $\rm{Tm^{3+}}$, but with opposite signs.
@@ Acknowledgements.
We acknowledge Andrew Wildes (ILL) for help with crystal orientation using OrientExpress. This work was supported by the DFG grant #SA 3688/1-1.

### v5

@@ V $\Gamma 4$ $A_{x}$
Figure 3: Schematic of the magnetic phase diagram for $\rm{TmFeO_{3}}$ a) external magnetic field along the b-axis b) along c-axis. Squares, circles and triangles are the inflection points of the temperature dependence function of the intensity for different reflections. Roman numbers designate regions corresponding by different magnetic representation: I – $\Gamma 4$; II – $\Gamma 24$; III – $\Gamma 2$; IV – $\Gamma^{\rm{Fe}}_{4}+\Gamma^{\rm{Tm}}_{2}$; V – $\Gamma^{\rm{Fe}}_{4}-\Gamma^{\rm{Tm}}_{4}$. Magnetic configurations, corresponding each region, are shown schematically without magnetic moments scaling.
@@
Regarding temperature region 2 – 30 K and field range 2 – 4 T, one can suppose, that in this case external magnetic field favors for the decoupling between $\rm{Tm^{3+}}$ and $\rm{Fe^{3+}}$ ions and Fe subsystem recover high-temperature configuration $\Gamma 4$, while Tm subsystem remains in $\Gamma 2$. At higher fields like 5 – 5.8 T Tm subsystem also goes to $\Gamma 4$ configuration but with ferromagnetic component directed opposite to $\rm{Fe^{3+}}$ one. The change of peaks intensity at magnetic field applied along b-axis and the consequent refinement of the magnetic structure confirm supposed phase transitions. Authors of [44] observed a magnetic-field-induced spin reorientation $\Gamma 2$ $\rightarrow$ $\Gamma 4$ at 1.6 K in a b-cut $\rm{TmFeO_{3}}$ single crystal in the static magnetic field range 2.2 – 3.6 T applied along b-axis using terahertz spectroscopy. We have also observed change of intensity of peaks at magnetic field applied along b-axis in the close field range that confirms phase transition. Situation with reverse reorientation transition without external magnetic field was observed in $\rm{TbFeO_{3}}$, a unique compound in which two types of spin reorientation from $\Gamma 4$ to $\Gamma 2$ at 8.5 K and then $\Gamma 2$ to $\Gamma 4$ at 3 K were reported [45].
Figure 4: Hysteresis of G-type peak (3,1,0-3,1,0) intensity in magnetic fields along the b-axis at 20 K. Error bars are of the size of the symbols.
In addition, hysteresis of the peak’s intensity for the field along the b-axis was observed in the experiment (Fig. 4), which was absent with the field along the c-axis, i.e. along the giant magnetocaloric axis. The largest difference between the peak intensities measured when the fields were increased and decreased was observed for G-type peaks like (3,1,0)(-3,1,0), presented at Figure 4, as it is provided by the strongest G-type ordered component of the magnetic structure. A smaller hysteresis was observed for A-type and F-type peaks, that connected with the smaller magnetic peaks’ intensities. Apparently, this difference is related with domains transformation when going from magnetic phase IV to phase V and back, which phases emerge when magnetic field applied along the b-axis. The domain structure in $\rm{TmFeO_{3}}$ was confirmed in work [32] where temperature magnetization hysteresis was described using model with a few domains, which explain jump/drop of magnetization in the spin reorientation temperatures range.
@@ IV Conclusions
X-ray studies demonstrate that the lattice parameters and atomic coordinates change almost linearly in $\rm{TmFeO_{3}}$ through the spin reorientation transition, keeping the orthorhombic symmetry group Pnma. The direction of the electric dipole moments does not change at any magnetic phase. This is additional evidence that electric system of $\rm{TmFeO_{3}}$ does not have influence on the spin-reorientation transition in this compound. The main reason for the spin reorientation transition connected with superexchange interactions 4f-3d, or with more complicate scheme of magnetic interactions [40]. Application of external magnetic field influences significantly on the spin-reorientation processes, moving the temperatures of the transitions up or low, depending on the direction of the external field. Thus, the application of the field along the c axis creates a situation favorable for the existence of the $\Gamma 2$ configuration, which leads to an increase in temperature of transition from $\Gamma 4$ to $\Gamma 2$ phase. The field along b-axis of $\rm{TmFeO_{3}}$ favors for $\Gamma 4$ configuration. This results in a decrease in the transition temperature from the $\Gamma 4$ phase to the $\Gamma 2$ phase, an increase in the temperature range in which the mixed phase is observed; and also allows for the formation of new phases. At low temperatures below 20 K and fields between 2 – 4 T a new magnetic configuration was observed, which consists of representation $\Gamma 4$ for $\rm{Fe^{3+}}$ ions and $\Gamma 2$ for $\rm{Tm^{3+}}$. At higher fields at these low temperatures another magnetic configuration emerges with representation $\Gamma 4$ for both $\rm{Fe^{3+}}$ subsystem and $\Gamma 4$ for $\rm{Tm^{3+}}$, but with opposite signs for $F_{z}$-components.
@@ Acknowledgements.
We acknowledge Andrew Wildes (ILL) for help with crystal orientation using OrientExpress. This work was supported by the DFG grant #SA 3688/1-1. The synthesis of single crystal $\rm{TmFeO_{3}}$ was carried out within the State assignment of the Ministry of Science and Higher Education of the Russian Federation for the Kirensky Institute of Physics, Krasnoyarsk Scientific Center, Siberian Branch of the Russian Academy of Sciences.